OpenStudy (jjuden):

What is the area of a circle whose radius is 20 meters?

OpenStudy (btaylor):

what is the formula for a circle's area?

OpenStudy (jjuden):

a 20pi m^2 b 40pi m^2 c 200pi m^2 d 400pim^2

OpenStudy (btaylor):

@jjuden in order to solve this you need to know the formula for a circle's area:\[A=\pi r^2\]What do you get when you plug in 20 for r?

OpenStudy (jjuden):

yes @BTaylor i already tried i did not get the answer

OpenStudy (btaylor):

If I were to rewrite it as \[A = r \times r \times \pi\] would it make more sense?

OpenStudy (jjuden):

i know how to do all that but im not getting what i need

OpenStudy (jjuden):

this is what i get 1256.64 square meters

OpenStudy (anonymous):

I don't really understand, it should be 1256. but your options are diffrent..

OpenStudy (btaylor):

You are going too far and evaluating. Each of your answers has something times pi, right?

OpenStudy (jjuden):

ikr @Light&Happiness

OpenStudy (btaylor):

What times pi = 1256.64?

OpenStudy (jjuden):

ok but where the heck does the m come from is what i wanna know

OpenStudy (btaylor):

m is for meters. It is the unit they give in the problem.

OpenStudy (jjuden):

ima just put d its the closest

OpenStudy (anonymous):

D is correct :o

OpenStudy (jjuden):

hey i got it right :D

OpenStudy (btaylor):

20 x 20 x pi = 400pi.

OpenStudy (jjuden):

it was D

OpenStudy (anonymous):

because 3.14 * 400 is 1256

OpenStudy (jjuden):

i got the D lol